### v1

# CARE: Confidence-Aware Regression Estimation of building density fine-tuning EO Foundation Models

Abstract
Performing accurate confidence quantification and assessment is important for deep neural networks to predict their failures, improve their performance and enhance their capabilities in real-world applications, for their practical deployment. For pixel-wise regression tasks, confidence quantification and assessment has not been well addressed in the literature, in contrast to classification tasks like semantic segmentation. The softmax output layer is not used in deep neural networks that solve pixel-wise regression problems. In this paper, to address these problems, we develop, train and evaluate the proposed model Confidence-Aware Regression Estimation (CARE). Our model CARE computes and assigns confidence to regression results. We focus on solving regression problems as downstream tasks of a Foundation Model for Earth Observation (EO). We evaluate the proposed model CARE and experimental results on data from the Copernicus Sentinel-2 satellite constellation for estimating the density of buildings show that the proposed method can be successfully applied to regression problems. We also show that our approach outperforms other methods.
Index Terms:
Earth Observation (EO), Foundation Model.
I Introduction
The utility of satellite data seems inexhaustible and, thanks to developments in Artificial Intelligence (AI), applications have begun to emerge at an accelerated pace. The focus is on developing methodological improvements in AI that: i) are needed for further improving real-world practicable solutions that use Earth Observation (EO) data, and ii) are prerequisites of important goals, namely accurately performing tasks that are “downstream” from EO data. The downstream tasks are, for example, building density estimation, land cover classification, crop type mapping and useful insights for Earth actions for sustainability and climate. Learning from unlabeled data, i.e. devising new methods for self-supervised learning, and developing an EO Foundation Model (FM) and evaluation framework with confidence quantification are important.
Self-supervised learning and operating within an EO FM framework are motivated by the lack of labeled data and the technical impossibility to provide labels for the massive volume of data collected by the European Space Agency (ESA) and other institutions. The trove of EO data collected is vast and expanding fast. The Sentinel-2 constellation generates approximately 1.6TB of compressed data daily. The lack of labels stands in the way of accomplishing important downstream tasks that we want to accurately perform. Such tasks that are “downstream” from EO data require the acquisition of labels. Also, because Earth is dynamic, labels change over time.
To use the large amounts of unlabeled data from satellites and achieve label efficiency, i.e. ability to learn from only few labeled data, EO FMs first perform self-supervised learning to extract general representations from unlabeled data and, then, supervised learning to specialize on specific applications, i.e. downstream tasks, by learning from a limited number of labeled data [1, 2]. For FMs, approximately $20\%$ of the labels for the use cases are needed, that would otherwise be required.
For deep neural networks that solve regression problems, confidence estimation is very different from and more challenging than for neural network models that solve classification problems. In addition, many more methods for confidence estimation have been proposed for classification neural networks, rather than for regression neural networks. The problem of accurately assigning a confidence metric to the real-valued output of a regression neural network is challenging and, to the best of the authors’ knowledge, has not been addressed well in the literature. Because the softmax output layer is not available in deep neural networks that solve regression problems, unlike in classification neural networks [4], the interpretation and the definition and meaning of confidence become less clear [6]. To address these problems, we propose a method that computes and assigns a confidence metric to the outputs of regression neural networks. We develop and train the proposed model Confidence-Aware Regression Estimation (CARE). Confidence is a proxy for the probability of correct classification. It is a metric between 0 and 1 that is an indicator of how well do we trust the results of the model. Confidence is an a priori estimate of the performance of our model.
Our main contributions are the development, training and evaluation of the proposed model CARE. We propose a confidence-aware inference method that first sorts the samples in the mini-batch with respect to their errors. In particular, we perform sorting of samples in ascending order, i.e. samples that have low errors go first in the mini-batch, while high errors go last. In this way, the confidence metric takes a value that represents the relative error level [6]. The model is then trained using this proposed mini-batch based training algorithm, where the outputs of the model CARE are both the confidence metric and the regression model output prediction. The desired output for the confidence metric result is based on sorting the samples in the mini-batches based on their error, i.e. absolute error or Mean Squared Error (MSE). Results on multi-spectral Sentinel-2 data, and more specifically for the problem of building density estimation, which is the regression task of predicting how close buildings are to each other, how densely built is the area in the image and how cities expand (i.e. urban growth), have shown that our method is effective.
II Related Work
Many different models exist that perform estimation and confidence assignment. However, most of the existing models are for classification (rather than regression) and use the softmax output probability [6]. Without model retraining, it is possible to correct, as well as calibrate to achieve no overfitting and use the entire range of possible output values, the output softmax probability of deep neural networks [4]. For example, by using temperature scaling, calibration based on the entropy and the difference between the first and second top class probabilities, or correction based on the aggregation of dispersion measures of softmax output probabilities over segments. Moreover, with model retraining, two-headed networks can be used to estimate the True Class Probability (TCP), which is a better estimator of confidence than Maximum Class Probability (MCP) [9].
For regression tasks, estimating model failures and regions of operation where the model should not be trusted is challenging [17]. We note that if we choose to turn a regression task to a classification problem, we lose precision, exactness, the desired outcome of having as an output a real-valued number, and having a result with an infinite number of decimal places. Considering the importance of solving regression problems, for example in EO for canopy height Above Ground Biomass (AGB) estimation or for CO2 flux estimation, it is crucial to develop methods that address the shortcomings of the existing methods and correctly compute and assign a confidence metric to the prediction outputs of regression neural networks [7].
Furthermore, most existing evaluation methods for EO FMs are based only on accuracy. However, EO datasets might have noisy labels and the evaluation of the performance of EO FMs using only the accuracy is not enough and might lead to misleading conclusions. Assessing the confidence with which the model performs inference is essential for usability and uptake. Accuracy, for the models and problems we consider, is not a sufficient performance measure. An evaluation framework for FMs has to include confidence assessment to benchmark and compare different FMs, evaluating their performance on both regression and classification segmentation downstream tasks.
III Our proposed methodology CARE
To increase reliability and trust, accurately computing the confidence with which the model performs inference and outputs its results is desirable for neural networks that solve regression problems. For models to function and be operational in the real-world, estimating and assigning an accurate confidence metric to every output model prediction is important. To successfully improve the performance of the model, assigning a confidence metric to every inference of the model is crucial, as for the correctly estimated low-confidence samples, the model can subsequently improve its performance on them. When we have samples that have low confidence, we need to perform further model retraining on these samples or collect more data near these samples, in this region in the data space. Using the model CARE, we are able to achieve improved performance by assigning a confidence metric to the output results, and this is one of the main reasons that confidence is important. Also, assigning a confidence metric to every output model prediction is essential, because for the low confidence data, the model can choose to abstain from providing a prediction in these specific cases, rather than outputting an incorrect estimation result.
The confidence metric helps humans to make correct decisions for critical matters, e.g. for policy. Moreover, to effectively combine possibly contradictory results from different models [6], the assigned confidence metrics can be used to better inform the fusion and integration of the different model outputs. Within a Mixture of Experts (MoE) framework, we can effectively combine different models using the confidence metric. The several different models might have outputs that lead to different findings and conclusions. In this way, we are able to associate a confidence measure to the inference of each of the models in the MoE and, then, combine their outputs to the final model decision using the individual confidences.
Using the confidence metric that we have assigned to the model output, we are also able to perform anomaly detection, as low confidence is an indicator of anomalies. This is important as anomaly detection/ Out-of-Distribution (OoD) detection is challenging, especially when the abnormalities are close to the normal/ in-distribution samples in the image data space [15, 16]. Confidence also enables us to perform confidence-based change detection, which is crucial in scenarios when we do not have labels for the change, as well as when the change is rare, occurring a small percentage of the time ($\leq 5\%$).
Loss function of CARE. The function minimized during training with Stochastic Gradient Descent (SGD) is given by
$\text{argmin}_{\boldsymbol{\theta}_{nn}}\,L(\textbf{x},\boldsymbol{\theta}_{nn},y,c,y^{*},c^{*})\text{,}$ (1)
where $\text{where }L=L_{0}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*})+\lambda\,L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,c,y^{*},c^{*})\text{,}$ (2)
where we denote the neural network parameters of our model by $\boldsymbol{\theta}_{nn}$. Here, the input image is $\textbf{x}_{i}$, the regression output $y_{i}$, the confidence output $c_{i}$, the ground truth regression value $y_{i}^{*}$ and the ground truth confidence $c_{i}^{*}$. The first loss term, $L_{0}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*})$, is the distance measure that is given by
$L_{0}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*})=\dfrac{1}{N}\sum_{i=1}^{N}||y_{i}-y_{i}^{*}||_{2}^{2}\text{,}$ (3)
where we compute the MSE and use the batch size $N$. The second loss term, $L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,c,y^{*},c^{*})$, is for the estimated confidence, where $\forall i,j$: $c_{i}\geq c_{j}\Leftrightarrow d(y_{i},y_{i}^{*})\leq d(y_{j},y_{j}^{*})$. This second loss term is the similarity metric that is given by
$L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*},c,c^{*})=\dfrac{1}{N}\sum_{i=1}^{N}|y_{i}-y_{i}^{*}|\times||c_{i}-c_{i}^{*}||_{2}^{2}\text{.}$ (4)
Architecture. The proposed model CARE estimates the confidence of the outputs of deep neural networks that solve pixel-wise regression problems. It assigns the confidence of $1$ for low error, and 00 for high error. Using the model CARE, we are able to estimate confidence as an unbiased predictor of the error rate. Here, for the implementation of the proposed methodology, we use the EO FM that we have recently trained, PhilEO [1, 3]. Within a FM framework, we are able to tackle and solve jointly a group of problems (downstream tasks) that are of interest to the EO community [5], [2]. For the training of the EO FM with unlabeled data, we have performed self-supervised learning that learns the correlations between: the Sentinel-2 multi-spectral images, masking and reconstruction, geo-location (longitude and latitude) estimation and climate zone classes. The architecture is a modified U-Net model [1]. For the downstream task of building density estimation, we fine-tune the EO FM, regress how close the buildings are in the image and evaluate the performance of the final model.
TABLE I: Evaluation of the proposed model CARE ($n=5000$), as well as comparison and ablation study, on the Sentinel-2 dataset PhilEO Bench Building Density Estimation (global, $10$ bands).
Model Error, Mean Error, Med. MSE MSE, $20\%$ MSE, $10\%$
CARE (Ours) 0.00759 0.00184 0.00326 $0.00057$ $0.00033$
Gaussian-Output NLL $0.00997$ 0.00123 $0.00468$ $0.00053$ $0.00036$
Error Sorting [6] $0.10586$ $0.03407$ 0.00334 0.00025 0.00025
Absolute Error model $0.00847$ $0.00254$ $0.00329$ $0.00056$ $0.00032$
The variation in the regression is modelled and captured by the confidence metric of CARE. In this paper, we focus on the task of building density regression. We have trained EO FMs using self-supervised learning, followed by supervised learning, making efficient use of labels [5], [1]. A methodology has been devised to take into account the geo-location longitude and latitude satellite information. As a downstream task, we retrain the model to regress building density. To accurately perform confidence estimation, the network has two output heads, i.e. one for the predicted regression and one for the estimated confidence, in (3) and (4). The latter is a metric indicating the extent to which we can trust the output results of the other head, i.e. predicted regression values [6, 7]. For $c_{i}^{*}$, in every mini-batch, after sorting the samples with respect to the error (i.e. MSE), $\eta=80\%$ of the data are assigned to $c_{i}^{*}=1$, while the remaining to $c_{i}^{*}=0$ [6]. This is a ranking problem, where correct predictions are ranked before incorrect ones. For the hyperparameter $\eta$, the threshold $80\%$ is chosen because we perform model retraining and first set $\lambda=0$ in (2) and, then, set $\lambda>0$ [9], [6]. The results of CARE in Sec. IV, because of the first term in (4), are not sensitive to relatively small changes in $\eta$. During inference, for our decision rule, to predict model failure, when the metric $(1-\text{confidence})$ is more than $\zeta=20\%$ of the regression value result, then we have detected incorrect pixels and CARE chooses to abstain from providing an output rather than predicting incorrect values.
IV Evaluation and Results of CARE
The evaluation of the proposed model CARE is based on the observation that the absolute error of the regression output should be equal to the quantity: $(1-\text{confidence})$. This should be true for every pixel and image. We examine if we learn to both: predict building density, and estimate confidence. We compute the average absolute error between: i) the absolute error of the regression output, and ii) the metric $(1-\text{confidence})$. The mean is over the pixels and the images, and the ideal result is zero. In the evaluation results, we are interested in the error between the regression error and the metric $(1-\text{confidence})$ being very small. Therefore, for the evaluation of the proposed model CARE and the estimated confidence, we use the Mean Absolute Error (MAE), and we also calculate the square Root MSE (RMSE). Furthermore, we compute the median absolute error, i.e. using the median instead of the average, where the median is more robust to outliers than the mean. The median error, in our case, is much lower than the average error.
TABLE II: Evaluation of the proposed model CARE on the dataset PhilEO Bench Building Density Estimation (Sentinel-2 data, L2A).
     $n$= 10000 7500 5000 1000 500 100 50
Error, Mean 0.00683 0.00761 0.00759 0.0138 $0.0167$ $0.0246$ $0.0261$
Error, Med. $0.00150$ $0.00190$ $0.00184$ $0.0065$ 0.0088 0.0147 0.0159
MSE $0.00301$ $0.00316$ $0.00326$ $0.0039$ 0.0043 0.0051 0.0034
MSE, $20\%$ $0.00062$ $0.00054$ $0.00057$ $0.0005$ 0.0006 0.0011 0.0011
MSE, $10\%$ $0.00031$ $0.00029$ $0.00033$ $0.0004$ 0.0005 0.0010 0.0007
In Table I, we evaluate CARE and examine both the average error and the median error. The median error of CARE is low, $0.00184$; this is desirable as it shows that the difference between the absolute error of the regression output and the estimated metric $(1-\text{confidence})$ is small. We also evaluate the regression output result using the MSE, like in (3), in Table I. We run experiments using $n$-sample tests, similar to [1], where we use $n$ training samples per region for fine-tuning. Stratified sampling is performed, as in the PhilEO Bench, and the regions are: Denmark, East Africa, Egypt, Guinea, Europe, Ghana, Israel, Japan, Nigeria, North America, Senegal, South America, Tanzania and Uganda [1, 5]. The performance of CARE, for $n=5000$, is examined in Table I.
For the evaluation of CARE, we also compare our model with other models. In Table I, we compare the results we obtain when we use Negative Log Likelihood (NLL) [17] and the model Gaussian-Output NLL. The percentage improvement of CARE compared to Gaussian-Output NLL is $30.34\%$ for the MSE, and $23.87\%$ for the average error. Moreover, we also compare CARE with the model Error-Sorting Confidence [6] in Table I. The percentage improvement of our model compared to Error-Sorting Confidence is $2.40\%$ for the MSE, and $92.83\%$ for the mean error. The percentage improvement of CARE compared to the model Absolute-Error Confidence is $0.91\%$ for the MSE, and $10.39\%$ for the average error. The Absolute-Error Confidence is also an ablation study, using only the first term in (4). These results demonstrate the efficacy of our algorithm and the superiority of CARE in outputting a reliable confidence metric for pixel-wise regression tasks, leading to both trustworthy and accurate model predictions.
TABLE III: Evaluation of the model Error-Sorting Confidence [6] on the PhilEO Bench building density regression, for comparison with the results of CARE in Table II, where Error Sorting is also an ablation study, i.e. using only the second term in (4).
     n= 10000 7500 5000 1000 500 100 50
Error, Mean 0.09969 0.08420 0.10586 0.0861 $0.0981$ $0.0963$ $0.1316$
Error, Med. $0.02332$ $0.02185$ $0.03407$ $0.0323$ 0.0417 0.0554 0.0765
MSE $0.00307$ $0.00320$ $0.00334$ $0.0039$ 0.0042 0.0056 0.0035
MSE, $20\%$ $0.00024$ $0.00027$ $0.00025$ $0.0004$ 0.0005 0.0013 0.0004
MSE, $10\%$ $0.00024$ $0.00027$ $0.00025$ $0.0004$ 0.0005 0.0013 0.0004
a) Input image
b) Prediction
c) Ground truth
d) Confidence
  map, CARE (Ours)
e) Abs. error btw
prediction & gt
f) Abs. error btw
pred. uncertainty & (e)
Figure 1: Pixel-wise regression and confidence estimation and assessment by the model CARE described in Sec. III on Sentinel-2 data.
We thus have compared CARE to the Gaussian-Output NLL model in Table I, where the latter minimizes the Gaussian negative logarithmic likelihood loss. Hence, the MSE loss for the regression error achieves higher accuracy (i.e. $0.00326$ compared to $0.00468$), and this is one of our main targets.
For the evaluation of CARE, in addition to the average and median absolute errors, we also compute the correlation between the confidence and the error in the regression. The correlation is $0.62090$. For the model Gaussian-Output NLL, it is $0.56696$. Here, the percentage improvement is $9.51\%$.
We also compare CARE to the ensembles method with $M=\{3,1\}$ members [18]. By combining the Gaussian distributions from the ensemble members using a heteroscedastic Gaussian log-likelihood loss, this method estimates the epistemic uncertainty. For $n=1000$, for $M=3$, the MSE and mean error are $0.00685$ and $0.01306$, respectively, while for $M=1$, these are $0.00696$ and $0.01353$. For CARE, in Table II, these are $0.0039$ and $0.0138$. Here, because of (2) and (3), the percentage improvement in MSE of our model CARE, compared to the $M=3$ ensembles method [18], is $43.1\%$.
Furthermore, for the evaluation of CARE, we also examine in Table I the MSE $20\%$ or $10\%$ which is a threshold, i.e. $\zeta$, for detecting instances where the model simply does not know the correct result (i.e. pixel-wise regression value) from the available input data, e.g. due to lack of spectral information or resolution. In such cases, models might choose to abstain from providing an answer and should be able to output “None of the above” for the segmentation regression result value. A building density of $20\%$ with an estimated error of ±$\pm 10\%$ leads to the predicted density of $20\%$ being not useful, i.e. not accurate. On the contrary, a building density of $80\%$ with an estimated error of ±$\pm 1\%$ leads to the predicted density of $80\%$ being useful and accurate. A prediction by the model with a high confidence indicates high reliability and trust for this particular prediction. Also, the features for high building density might not be clear in multi-spectral optical EO data. This is due to epistemic uncertainty, as it is induced by the lack of detail in the measurement. This is why a plausible output set that is an Open Set, is needed. Epistemic uncertainty is systematic, caused by lack of knowledge, and can be reduced by learning the characteristics of the quantity (e.g., high building density) using additional information (for example, in-situ measurements). Moreover, aleatoric uncertainty is statistical and related to randomness, and the sample not being a typical example of the quantity (e.g., high or low building density).
a) Input image
b) Prediction
c) Ground truth
d) Confidence
  map, CARE (Ours)
e) Abs. error btw
prediction & gt
f) Abs. error btw
pred. uncertainty & (e)
Figure 2: CARE regression and confidence estimation: Building density.
Our experiments in Tables I-III show that CARE is effective and outperforms other models. The focal point of this paper is confidence quantification and assessment for pixel-wise regression tasks in EO using neural networks that have a continuous output. CARE achieves good generalisation performance, and this work’s methodological and model development value, as well as its application implementation value, is high. The obtained results can be useful for researchers and practitioners (i.e. from a practical viewpoint). Using a mathematical definition for the confidence metric for pixel-wise regression tasks, a contribution of this paper is the value for applications, so that researchers studying such real-world problems can take advantage of the results and the attained good performance.
We evaluate the proposed model CARE in Table II using $n$-sample tests, for $n$ from $50$ to $10000$ samples. We examine both the MSE for the accuracy of the regression, i.e. (3) in (2), and the average error for the correctness of the confidence metric, i.e. of $c_{i}$ in (4). Here, we observe that for $n=7500$ samples, the MSE is $0.00316$, while for $n=500$, the MSE is $0.0043$, i.e. the performance of the model improves when we have more labeled data. In addition, for the accuracy of the confidence metric of our model, the median error in Table II is $0.00190$ for $n=7500$ samples, while the median error is $0.0088$ for $n=500$. The performance of the model CARE, for confidence quantification, improves when we have more labeled data, i.e. $n=7500$ samples compared to $n=500$.
For $n=1000$ in Table II, CARE achieves the MSE of $0.0039$, for the accuracy of the building density regression. The model Gaussian-Output NLL from Table I, for $n=1000$, yields the MSE of $0.0052$. Here, the percentage improvement of our model, compared to Gaussian-Output NLL, is $25\%$.
For the evaluation of our model CARE, we also examine in Table II the MSE $20\%$ or $10\%$, similar to in Table I. We observe that the performance of the proposed model CARE for estimating the building density, i.e. how close the buildings are in the Sentinel-2 multi-spectral images, improves when the model has the ability to abstain in cases of low confidence. These results demonstrate the effectiveness of our algorithm in providing a reliable confidence metric for regression that leads to trustworthy and more accurate model predictions.
a) Input image
b) Prediction
c) Ground truth
d) Confidence
  map, CARE (Ours)
e) Abs. error btw
prediction & gt
f) Abs. error btw
pred. uncertainty & (e)
Figure 3: Regression and confidence quantification: CARE using (1)-(4).
In Table III, we evaluate the model Error-Sorting Confidence [6] on the Sentinel-2 PhilEO Bench building density estimation downstream task [1], to compare these results with the results of the proposed model CARE in Table II. Error-Sorting Confidence is used in the fine-tuning stage, i.e. when starting from and using a trained EO FM [1, 3]. We observe that for $n=100$ training samples per region, the percentage improvement of CARE compared to Error-Sorting Confidence is $8.93\%$ for the MSE, and $73.47\%$ for the median error.
For the evaluation of CARE, we also examine qualitative results, in Figs. 1-3. For the error in the prediction, i.e. in (e), the variation in the regression output is, for example, because we underestimate in (b) building density. In (f), the predicted uncertainty is $(1-\text{confidence})$, i.e. using (d). In the colorbar in (f), the maximum value is smaller than $1$, and this is desirable, i.e. we have less highlighted and emphasized yellow color.
V Conclusion
In this paper, we have proposed the model CARE for pixel-wise regression tasks. Our model CARE computes and assigns a confidence metric to the output results of deep neural networks that solve regression problems. The focal point of this work has been confidence quantification and assessment of regression algorithms that are based on EO Foundation Models applied on downstream tasks, i.e. building density estimation from satellite Sentinel-2 data. Using the proposed model CARE, we are able to estimate confidence as an unbiased predictor of the error rate. In this work, we have evaluated the proposed model CARE using the average absolute error between the absolute error of the regression output and the metric $(1-\text{confidence})$. We have demonstrated that the model CARE is effective and is able to both predict building density and estimate confidence. In addition, we have also compared our model with other baseline models and we have shown that the proposed model CARE outperforms other approaches.