### v2

# CARE: Confidence-Aware Regression Estimation of building density fine-tuning EO Foundation Models

Abstract
Performing accurate confidence quantification and assessment in pixel-wise regression tasks, which are downstream applications of AI Foundation Models for Earth Observation (EO), is important for deep neural networks to predict their failures, improve their performance and enhance their capabilities in real-world applications, for their practical deployment. For pixel-wise regression tasks, specifically utilizing remote sensing data from satellite imagery in EO Foundation Models, confidence quantification is a critical challenge. The focus of this research is on developing a Foundation Model using EO satellite data that computes and assigns a confidence metric alongside regression outputs to improve the reliability and interpretability of predictions generated by deep neural networks. To this end, we develop, train and evaluate the proposed Confidence-Aware Regression Estimation (CARE) Foundation Model. Our model CARE computes and assigns confidence to regression results as downstream tasks of a Foundation Model for EO data, and performs a confidence-aware self-corrective learning method for the low-confidence regions. We evaluate the model CARE, and experimental results on multi-spectral data from the Copernicus Sentinel-2 constellation to estimate the building density (i.e. monitoring urban growth), show that the proposed method can be successfully applied to important regression problems in EO. We also show that our model CARE outperforms other methods.
Index Terms:
Earth Observation (EO), Pixel-wise regression, EO Foundation Models, Remote sensing, Copernicus Sentinel-2
I Introduction
The significance of confidence quantification and assessment in deep learning, specifically in AI Foundation Models in Earth Observation (EO) that use satellite data, for regression applications is critical. The utility of satellite data seems inexhaustible, and thanks to developments in AI, applications emerge at an accelerated pace in EO Foundation Models using remote sensing data. The trove of EO data is vast and expanding fast. The ESA Sentinel-2 constellation generates approximately 1.6TB of compressed multi-spectral data daily. However, the lack of labels stands in the way of accurately performing important downstream tasks, and confidence quantification is critical to achieve better performance.
The focus of this work is on developing methodological improvements in EO Foundation Models that: i) are needed to further improve performance for practicable real-world solutions that use EO data, and ii) are prerequisites to achieve important goals, namely accurately performing tasks that are ‘downstream’ from remote sensing data. The downstream tasks are, for example, building density estimation, land cover classification, crop type mapping and useful insights for Earth actions for sustainability and climate. Learning from unlabeled data, i.e. devising new methods for self-supervised learning, and developing an EO Foundation Model (FM) with confidence quantification and assessment are important. Confidence is a proxy for the probability of correct results. It is a metric between 00 and $1$ that is an indicator of how well do we trust the output results of the model. Confidence is an a priori estimate of the performance of our model.
The problem of accurately assigning a confidence metric to the real-valued output of a pixel-wise regression neural network in EO Foundation Models to estimate building density from satellite data is challenging and, to the best knowledge of the authors, has not been well addressed in the literature.
Our main contributions. To address these problems, we propose the Confidence-Aware Regression Estimation (CARE) Foundation Model for remote sensing and EO. The model CARE computes and assigns confidence metrics alongside regression outputs in order to improve the reliability and interpretability of predictions generated by deep neural networks. The CARE model is implemented based on a modified U-Net architecture suitable for processing multi-spectral satellite images, which allows it to handle the complex spatial features present in the data. Assigning a confidence metric to every inference of the model is crucial, as for the correctly estimated low-confidence samples in building density estimation from satellite data, our model subsequently implements a self-corrective method and improves its performance on them. CARE uses two heads that compute and assign a regression density measure and a confidence metric to the output of regression neural networks. Using our proposed loss function that comprises a Mean Squared Error (MSE) distance loss and a confidence loss, our model successfully performs self-corrective confidence-aware regression for building density estimation which is a downstream task of EO Foundation Models, and outperforms other methods.
Our main contributions are the development, training and evaluation of the proposed CARE model. Results on multi-spectral Sentinel-2 data, and more specifically for the problem of building density estimation, which is the regression task of predicting how close buildings are to each other, how densely built is the area in the image and how cities expand (i.e. urban growth) using EO Foundation Models, have shown that our method is effective and outperforms other base models.
II Related Work
Foundation Models in EO. To use the vast amounts of unlabeled satellite data and achieve label efficiency, i.e. the ability to learn from only few labeled data in fine-tuning, EO Foundation Models are first pretrained on large unlabeled satellite datasets by performing self-supervised learning to extract special representations from unlabeled data [2, 33]. Then, in the downstream applications, Foundation Models are fine-tuned by performing supervised learning to specialize in specific applications, i.e. downstream tasks, by learning from a limited number of labeled data. In inference, classification or regression is performed for the downstream task [1, 2].
Self-supervised learning and operating within an EO FM framework [24, 23] are motivated by the lack of labeled data and the technical impossibility to provide labels for the massive volume of data collected by the European Space Agency (ESA) and other space agencies (e.g., NASA, JAXA). Enhancing EO Foundation Models with confidence quantification and assessment is critical to improve their performance.
Confidence-aware models. There are many different models that perform estimation and confidence assignment [7, 26, 28]. However, most of the existing models are not for pixel-wise regression building density estimation. Existing models for classification are based on the softmax output probability [6]. For regression problems where the result is a real-valued number, the output softmax layer is not used in the neural network [21]. Therefore, for regression neural networks that do not use softmax, it is not straightforward to apply softmax-probability-based classification confidence methods.
Without model retraining, it is possible to correct, as well as calibrate to achieve no overfitting and use the entire range of possible output values, the output softmax probability of deep neural networks [4]. For example, by using temperature scaling: Calibration based on entropy and the difference between the first and second top class probabilities, or correction based on the aggregation of dispersion measures of softmax output probabilities over segments. Moreover, with model retraining, two-headed networks can be used to estimate the True Class Probability (TCP), which is a better confidence estimator than Maximum Class Probability (MCP) [9].
For regression tasks, estimating model failures and regions of operation where the model should not be trusted is challenging [17]. We note that if we choose to turn a regression task into a classification problem [26], then we lose precision, exactness and the desired outcome of having as an output a real-valued number.
Considering the importance of solving regression problems [21], for example in EO for the accurate estimation of canopy height Above Ground Biomass (AGB) [2], or for the estimation of CO2 flux, it is crucial to develop methods that address the shortcomings of existing methods and correctly compute and assign a confidence metric to the prediction outputs of regression neural networks [7].
The work in [19] proposes a spatial self-corrective learning framework which: a) is able to locate predictions with potential large errors by estimating confidence using Maximum Likelihood Estimation (MLE), b) uses confidence-based pseudo-interpolation to correct low-confidence predictions in local neighborhoods, and c) performs recurrent self-refinement of the estimated height in an iterative manner. Experiments on different landscapes in the high-latitude regions show improvements compared to the other baseline methods.
The work in [20] presents a probabilistic regression method. The deep neural network estimates the unnormalized density from the input-target pair using a conditional target density $p(y|x)$ model, with energy as the basis $(x,y)$. Monte Carlo sampling is used to reduce the negative log probability of this probabilistic $p(y|x)$ model.
Furthermore, in the literature [20, 29, 30, 31], in addition to probabilistic models and Bayesian neural networks, for confidence assignment and assessment in regression neural networks, evidential regression can also be performed [35, 36, 37], as well as conformal prediction (i.e. conformal inference) for regression [38, 39].
EO data noisy labels. Most existing evaluation methods for EO Foundation Models [2, 40] are based only on accuracy. However, remote sensing datasets might have noisy labels, i.e. incorrect annotations, and the evaluation of the performance of EO FMs using only the accuracy is not enough and might lead to misleading conclusions. Assessing the confidence with which the model performs inference is essential for usability and uptake [32, 25]. Accuracy, on its own, for the models and problems we consider, is not a sufficient performance measure. An evaluation framework for FMs has to include confidence assessment to benchmark and compare different Foundation Models, evaluating their performance on both regression and classification segmentation downstream tasks.
Confidence metric. The confidence metric helps humans to make correct decisions for critical matters [9, 27, 22], e.g. for policy. Moreover, to effectively combine possibly contradictory results from different models [6], the assigned confidence metrics can be used to better inform the integration and fusion of the different model outputs. Within a Mixture of Experts (MoE) framework, we can effectively combine different models using the confidence metric. Several different models might have outputs that lead to different findings and conclusions. In this way, we are able to associate a confidence measure to the inference of each of the models in the MoE and, then, combine their outputs to the final model decision using the individual confidences.
Using the confidence metric that we have assigned to the model output, we are also able to perform anomaly detection, as low confidence is an indicator of anomalies. This is important as anomaly detection/ Out-of-Distribution (OoD) detection is challenging, especially when the abnormalities are close to the normal/ in-distribution samples in the image data space [15, 16]. In addition, the confidence metric also enables us to perform confidence-based change detection, which is crucial in scenarios when we do not have labels for the change, as well as when the change is rare, occurring only a small percentage of the time (e.g., $<5\%$, or even $<2\%$).
Figure 1: CARE for confidence-aware regression: assign confidence metric to predictions, identify wrong predictions and refine the model.
III Our proposed methodology
Confidence-aware pixel-wise regression. To increase reliability and trust, accurately computing the confidence with which the model performs inference and outputs its results is desirable for neural networks that solve regression problems. For models to function and be operational in the real-world, estimating and assigning an accurate confidence metric to every output model prediction is important. To successfully improve the performance of the model, assigning a confidence metric to every inference output of the model is crucial, as for the correctly estimated low-confidence samples, the model can subsequently improve its performance specifically on them. When we have samples with low confidence, we need to perform further retraining of the model on these samples or collect more data near these samples (or data augmentation), in this region in the data space. Using the model CARE, we are able to achieve improved performance by assigning a confidence metric to the output results, and this is one of the main reasons that confidence is important. Also, assigning a confidence metric to every output model prediction is essential, because for the low confidence data, the model can choose to abstain from providing a prediction in these specific cases, rather than outputting an incorrect estimation result.
III-A Proposed method
The core methodology of the model CARE is based on a mini-batch training algorithm that takes advantage of the computed MSE distance measure assigned to the regression predictions. CARE uses two heads that compute and assign a regression building density measure and a confidence metric to the output of regression neural networks. The approach is built upon sorting samples within mini-batches based on their prediction absolute errors, which allows for establishing a relative confidence metric for each sample.
Sorting mechanism. In each mini-batch, samples are sorted in an ascending order of their errors (using the MSE as a measure of error). This enables the model to assign higher confidence scores to samples with lower errors and vice versa.
Training procedure with dual outputs. The model operates with two output heads: one for the regression prediction and another for the confidence metric. The training of the model aims to minimize a custom loss function that consists of two components: a) Regression Loss (i.e. $L_{0}$): Measures the distance between the predicted outputs and the ground truth values using the MSE. b) Confidence Loss ($L_{1}$): Assesses the relationship between the prediction absolute errors and the assigned confidence scores, ensuring that higher confidence correlates with lower prediction errors.
Loss function. The function minimized during training with Stochastic Gradient Descent (SGD) is given by:
$\text{argmin}_{\boldsymbol{\theta}_{nn}}\,L(\textbf{x},\boldsymbol{\theta}_{nn},y,c,y^{*},c^{*})\text{,}$ (1)
where $\text{where }L=L_{0}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*})+\lambda\,L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,c,y^{*},c^{*})\text{,}$ (2)
where we denote the neural network parameters of our model by $\boldsymbol{\theta}_{nn}$. In the combined loss function (1), $\lambda$ is a hyperparameter that can be adjusted to weigh the importance of confidence estimation relative to regression accuracy during training. Here, the input image is $\textbf{x}_{i}$, the regression output $y_{i}$, the confidence output $c_{i}$, the ground truth regression value $y_{i}^{*}$ and the ground truth confidence $c_{i}^{*}$. The first loss term, $L_{0}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*})$, is the distance measure that is given by:
$L_{0}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*})=\dfrac{1}{N}\sum_{i=1}^{N}||y_{i}-y_{i}^{*}||_{2}^{2}\text{,}$ (3)
where we compute the MSE and use the batch size $N$. The second loss term, $L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,c,y^{*},c^{*})$, is for the estimated confidence, where $\forall i,j$: $c_{i}\geq c_{j}\Leftrightarrow d(y_{i},y_{i}^{*})\leq d(y_{j},y_{j}^{*})$. This second loss term is the confidence metric that is given by:
$L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*},c,c^{*})=\dfrac{1}{N}\sum_{i=1}^{N}|y_{i}-y_{i}^{*}|\cdot||c_{i}-c_{i}^{*}||_{2}^{2}\text{.}$ (4)
Architecture. The model CARE is implemented based on a modified U-Net architecture that has approximately $70$M parameters, suitable for processing multi-spectral satellite images in EO Foundation Models, which allows it to handle the complex spatial features present in the data [1]. The U-Net’s encoder-decoder structure is particularly adept at tasks requiring precise localization, important for pixel-wise predictions in EO tasks.
Datasets. Our model CARE is pretrained on the PhilEO Globe Sentinel-2 global dataset [1, 3] and fine-tuned on the PhilEO Bench downstream tasks dataset [1]. CARE and the other examined baseline models in this work are evaluated on the PhilEO Bench [1, 3, 4, 5].
III-B CARE flowchart diagram
The proposed model CARE estimates the regression value and the confidence metric in the output of the deep neural network that solves pixel-wise regression problems in Fig. 1. CARE assigns the confidence of $1$ for low error, and 00 for high error.
For the training of the EO Foundation Model with unlabeled data, we have performed self-supervised learning that learns the correlations between: the Sentinel-2 multi-spectral images, masking and reconstruction, geolocation (longitude and latitude) estimation and climate zone classes. For the downstream task of building density estimation, we fine-tune the EO Foundation Model, regress how close the buildings are in the image and evaluate the performance of the final model.
In this paper, we focus on the task of building density regression. To accurately perform confidence estimation, the network has two output heads, i.e. one for the predicted regression output value and one for the estimated confidence, in (3) and (4). The latter is a confidence metric that correlates with the ground truth, indicating the extent to which we can trust the output results of the other head, i.e. the predicted regression values [6, 7].
For $c_{i}^{*}$, in every mini-batch, after sorting the samples in an ascending order with respect to their errors (i.e. MSE), $\eta=80\%$ of the data are assigned with high confidence scores (i.e. $c_{i}^{*}=1$), while the remaining data to low confidence scores (i.e. $c_{i}^{*}=0$). For the hyperparameter $\eta$, the threshold $80\%$ is chosen because we perform model retraining, and we first set $\lambda=0$ in (2) and, then, set $\lambda>0$ [9], [6]. During inference, for our decision rule, to predict model failure, when the metric $(1-\text{confidence})$ is more than $\zeta=20\%$ (or $10\%$) of the regression value result, then we have detected incorrect pixels and CARE performs self-correction by choosing to abstain from providing an output rather than predicting incorrect regression values.
TABLE I: Evaluation of the proposed model CARE ($n=5000$), as well as comparison and ablation study, on the Sentinel-2 dataset PhilEO Bench Building Density Estimation (global, $10$ bands).
Model Error, Mean Error, Med. MSE MSE, $20\%$ MSE, $10\%$
CARE (Ours) 0.00759 0.00184 0.00326 $0.00057$ $0.00033$
Gaussian-Output NLL $0.00997$ 0.00123 $0.00468$ $0.00053$ $0.00036$
Error Sorting [6] $0.10586$ $0.03407$ 0.00334 0.00025 0.00025
Absolute Error model $0.00847$ $0.00254$ $0.00329$ $0.00056$ $0.00032$
IV Evaluation and Results
The evaluation of the proposed model CARE is based on the observation that the absolute error of the regression output should be equal to the quantity: $(1-\text{confidence})$. This should be true for every pixel and image. We examine if we learn to both: predict building density, and estimate confidence. We compute the average absolute error between: i) the absolute error of the regression output, and ii) the metric $(1-\text{confidence})$. The mean is over the pixels and the images, and the ideal result is zero. In the evaluation results, we are interested in the error between the regression absolute error and the metric $(1-\text{confidence})$ being very small. Therefore, in this work, for the evaluation of the proposed model CARE and the estimated confidence, we use the MSE. Furthermore, we compute the median absolute error, i.e. using the median instead of the average, where the median is more robust to outliers than the mean. The median absolute error, in our case for building density estimation, is much lower than the average error.
IV-A Evaluation of CARE using absolute error in Table I and comparison to other models
In Table I, we evaluate the proposed model CARE and examine both the average error and the median error for the results of the model. The median error of the proposed model CARE is low, i.e. $0.00184$, and this is desirable as it shows that the difference between the absolute error of the regression output and the estimated metric $(1-\text{confidence})$ is small. We also evaluate the regression output result using the MSE, in Tables I and II. We run experiments using $n$-sample tests, where we use $n$ training samples per region for fine-tuning. Stratified sampling is performed, and the different regions are: Denmark, East Africa, Egypt, Guinea, Europe, Ghana, Israel, Japan, Nigeria, North America, Senegal, South America, Tanzania and Uganda [1, 5]. The performance of the model CARE, for $n=5000$, is examined in Table I.
TABLE II: Evaluation of the proposed model CARE on the dataset PhilEO Bench Building Density Estimation (Sentinel-2 data, L2A).
     $n$= 10000 7500 5000 1000 500 100 50
Error, Mean 0.00683 0.00761 0.00759 0.0138 $0.0167$ $0.0246$ $0.0261$
Error, Med. $0.00150$ $0.00190$ $0.00184$ $0.0065$ 0.0088 0.0147 0.0159
MSE $0.00301$ $0.00316$ $0.00326$ $0.0039$ 0.0043 0.0051 0.0034
MSE, $20\%$ $0.00062$ $0.00054$ $0.00057$ $0.0005$ 0.0006 0.0011 0.0011
MSE, $10\%$ $0.00031$ $0.00029$ $0.00033$ $0.0004$ 0.0005 0.0010 0.0007
For the evaluation of CARE, we also compare our model with other models. In Table I, we compare the results we obtain when we use Negative Log Likelihood (NLL) [17] and the model Gaussian-Output NLL. The percentage improvement of CARE compared to Gaussian-Output NLL is $30.34\%$ for the MSE, and $23.87\%$ for the average error. We observe that according to these results, for the use case of building density estimation, regarding the model Gaussian-Output NLL, an interesting open research question is how useful the normal distribution assumption is. Moreover, we also compare CARE with the model Error-Sorting Confidence [6] in Table I. The percentage improvement of our model compared to Error-Sorting Confidence is $2.40\%$ for the MSE, and $92.83\%$ for the mean absolute error.
The percentage improvement of CARE compared to the model Absolute-Error Confidence is $0.91\%$ for the MSE, and $10.39\%$ for the average error. The Absolute-Error Confidence is also an ablation study, using only the first term in (4). These results demonstrate the efficacy of our algorithm and the superiority of CARE in outputting a reliable confidence metric for pixel-wise regression tasks, leading to both trustworthy and accurate model predictions.
TABLE III: Evaluation of the model Error-Sorting Confidence [6] on the PhilEO Bench building density regression, for comparison with the results of CARE in Table II, where Error Sorting is also an ablation study, i.e. using only the second term in (4).
     n= 10000 7500 5000 1000 500 100 50
Error, Mean 0.09969 0.08420 0.10586 0.0861 $0.0981$ $0.0963$ $0.1316$
Error, Med. $0.02332$ $0.02185$ $0.03407$ $0.0323$ 0.0417 0.0554 0.0765
MSE $0.00307$ $0.00320$ $0.00334$ $0.0039$ 0.0042 0.0056 0.0035
MSE, $20\%$ $0.00024$ $0.00027$ $0.00025$ $0.0004$ 0.0005 0.0013 0.0004
MSE, $10\%$ $0.00024$ $0.00027$ $0.00025$ $0.0004$ 0.0005 0.0013 0.0004
We therefore have compared CARE to the Gaussian-Output NLL model in Table I, where the latter minimizes the Gaussian negative logarithmic likelihood loss. Hence, the MSE loss for the regression error achieves higher accuracy (i.e. $0.00326$ compared to $0.00468$), and this is one of our main targets. For the evaluation of CARE, in addition to the average and median absolute errors, we also compute the correlation between the confidence and the error in the regression. The correlation is $0.62090$. For the model Gaussian-Output NLL, it is $0.56696$. Here, the percentage improvement is $9.51\%$.
We also compare the proposed model CARE to the ensembles method with $M=\{3,1\}$ members [18]. By combining the Gaussian distributions from the ensemble members using a heteroscedastic Gaussian log-likelihood loss, this method estimates the epistemic uncertainty. For $n=1000$, for $M=3$, the MSE and mean absolute error are $0.00685$ and $0.01306$, respectively, while for $M=1$, these are $0.00696$ and $0.01353$. For CARE, in Table II, these are $0.0039$ and $0.0138$. Here, because of (2) and (3), the percentage improvement in MSE of our model CARE, in comparison to the $M=3$ ensembles method [18], is $43.1\%$.
IV-B Further evaluation of CARE based on threshold $\zeta$, Table I
Furthermore, for the evaluation of CARE, we also examine in Table I the MSE $20\%$ or $10\%$ which is a threshold, i.e. $\zeta$, to detect instances where the model simply does not know the correct result (i.e. pixel-wise regression value) from the available input data, e.g. due to lack of spectral information or resolution. In such cases, models might choose to abstain from providing an answer and should be able to output “None of the above” for the result value of segmentation. A building density of $20\%$ with an estimated error of ±$\pm 10\%$ leads to the predicted density of $20\%$ being not useful, i.e. not accurate. In contrast, a building density of $80\%$ with an estimated error of ±$\pm 1\%$ leads to the predicted density of $80\%$ being useful and accurate.
a) Input image
b) Prediction
c) Ground truth
d) Confidence
  map, CARE (Ours)
e) Abs. error btw
prediction & gt
f) Abs. error btw
pred. uncertainty & (e)
Figure 2: Pixel-wise regression and confidence estimation and assessment by the model CARE described in Sec. III on Sentinel-2 data.
A prediction of the model with a high confidence indicates high reliability and trust for this particular prediction. Also, the features for high building density might not be clear in multi-spectral optical EO data. This is due to epistemic uncertainty, as it is induced by the lack of detail in the measurement. This is why a plausible output set that is an Open Set, is needed. Epistemic uncertainty is systematic, caused by lack of knowledge, and can be reduced by learning the characteristics of the quantity (e.g., high building density) using additional information (for example, in-situ measurements). Moreover, aleatoric uncertainty is statistical and related to randomness, and the sample not being a typical example of the quantity (e.g., very high or low building density).
a) Input image
b) Prediction
c) Ground truth
d) Confidence
  map, CARE (Ours)
e) Abs. error btw
prediction & gt
f) Abs. error btw
pred. uncertainty & (e)
Figure 3: CARE regression and confidence estimation: Building density.
Our experiments in Tables I-III show that CARE is effective and outperforms other models. The focal point of this paper is confidence quantification and assessment for pixel-wise regression tasks in EO using neural networks that have a continuous output. CARE achieves good generalisation performance, and this work’s methodological and model development value, as well as its application implementation value, is high. The obtained results can be useful for researchers and practitioners (i.e. from a practical viewpoint). Using a mathematical definition for the confidence metric for pixel-wise regression tasks, a contribution of this paper is the value for applications, so that researchers studying such real-world problems can take advantage of the results and the attained good performance.
IV-C Evaluation of CARE using n-shot experiments in Table II
We evaluate the proposed model CARE in Table II using $n$-sample tests, for $n$ from $50$ to $10000$ samples. We examine both the MSE for the accuracy of the regression, i.e. (3) in (2), and the average error for the correctness of the confidence metric, i.e. $c_{i}$ in (4). Here, we observe that for $n=7500$ samples, the MSE is $0.00316$, while for $n=500$, the MSE is $0.0043$, i.e. the performance of the model improves when we have more labeled data. In addition, for the accuracy of the confidence metric of our model, the median error in Table II is $0.00190$ for $n=7500$ samples, while the median error is $0.0088$ for $n=500$. The performance of the model CARE, for confidence quantification, improves when we have more labeled data, i.e. $n=7500$ samples compared to $n=500$.
For $n=1000$ in Table II, CARE achieves the MSE of $0.0039$, for the accuracy of the building density regression. The model Gaussian-Output NLL from Table I, for $n=1000$, yields the MSE of $0.0052$. Here, the percentage improvement of our model, compared to Gaussian-Output NLL, is $25\%$.
For the evaluation of our model CARE, we also examine in Table II the MSE $20\%$ or $10\%$, similar to in Table I. We observe that the performance of the proposed model CARE for estimating the building density, i.e. how close the buildings are in the Sentinel-2 multi-spectral images, improves when the model has the ability to abstain in cases of low confidence. Moreover, Table II shows that when assuming that $10\%$ provide low confidence and our model should abstain, CARE gives improved performance, compared to the assumption of $20\%$. These results demonstrate the effectiveness of our algorithm in providing a reliable confidence metric for regression that leads to trustworthy and more accurate model predictions.
a) Input image
b) Prediction
c) Ground truth
d) Confidence
  map, CARE (Ours)
e) Abs. error btw
prediction & gt
f) Abs. error btw
pred. uncertainty & (e)
Figure 4: Regression and confidence quantification: CARE using (1)-(4).
IV-D Baseline model n-shot evaluation, Table III
In Table III, we evaluate the model Error-Sorting Confidence [6] on the Sentinel-2 PhilEO Bench building density estimation downstream task [1], to compare these results with the results of our proposed model CARE in Table II. Error-Sorting Confidence is used in Table III in the fine-tuning stage, i.e. when starting from and using a trained EO FM [1, 3]. Here, Error-Sorting is also an ablation study, i.e. using only the second term in (4). We observe that for $n=100$ training samples per region, the percentage improvement of CARE compared to Error-Sorting Confidence is $8.93\%$ for the MSE, and $73.47\%$ for the median error.
IV-E Qualitative evaluation of CARE
For the evaluation of CARE, we also examine qualitative results, in Figs. 2-4. In addition to the numerical evaluation metrics presented in the preceding subsections, we also evaluate the proposed model CARE qualitatively. For the error in the prediction, i.e. in (e), the variation in the regression output is, for example, because we underestimate in (b) the building density. In (f), the predicted uncertainty is $(1-\text{confidence})$, i.e. using (d). In the colorbar in (f), the maximum value is smaller than $1$, and this is desirable, i.e. we have less highlighted and emphasized yellow color.
V Conclusion
In this paper, we have proposed the model CARE for pixel-wise regression tasks, for building density estimation from satellite Sentinel-2 data. Our model CARE computes and assigns a regression measure and a confidence metric to the output results of deep neural networks that solve regression problems. CARE not only predicts building density effectively but also provides reliable confidence scores that indicate trust in its predictions. The results reveal that CARE achieves a percentage improvement over Gaussian-Output NLL and other confidence estimation models. Evaluation metrics include the MSE, the average absolute error and the median absolute error, allowing comprehensive performance analysis against several baseline models. The experiments demonstrate that CARE significantly outperforms other baseline methods, showing improvements in various metrics. CARE shows a notable reduction in prediction error and demonstrates a strong correlation between confidence values and prediction errors. The results summarize the innovation introduced by the proposed model CARE, underscoring its dual capability of providing both accurate regression outputs and meaningful confidence metrics. We have shown that confidence quantification and assessment is critical for the deployment of models in real-life applications, particularly in EO tasks where uncertainties can have significant implications. By addressing the gap in confidence estimation for regression tasks in EO Foundation Models for building density estimation, this work contributes a valuable methodology and insights that can further enhance the practical application of neural networks in geospatial analysis. This underscores the importance of developing models that not only provide predictions but also quantify certainty, facilitating informed decision-making in critical areas such as urban planning and environmental monitoring.